OpenStudy (anonymous):

what is the solution of this linear system 3x+y=-2 y=-x+4

OpenStudy (amistre64):

what does y equal?

OpenStudy (amistre64):

they tell you up front that: y=-x+4 use that to solve for x. All the y parts in the system have to be the same ....

OpenStudy (amistre64):

the first equation is setup as: 3x + y = -2 ^^ we know this is equal to -x+4 3x + (-x+4) = -2 now it is relatively simple to find a solid value for x

OpenStudy (amistre64):

when we have a known value for x, we can then use it to define y=-x+4

OpenStudy (anonymous):

3x - 1x + 4 = -2 2x +4= -2 2x = -6 x= -3

OpenStudy (amistre64):

good job :) since x = -3 y = -(-3) + 4

OpenStudy (anonymous):

y = 5

OpenStudy (amistre64):

not quite

OpenStudy (amistre64):

y = -(-3) + 4 y = 3 + 4

OpenStudy (anonymous):

i know i messed up on it lol y= 7

OpenStudy (amistre64):

then lets go with the (-3,7) ;) good job